• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ramesh

    Log entry time 15:59:33 on May 1,2011

    Entry number 357741

    keyword=Shift Summary

    08:00 Shift starts with no beam (since about 02:00) due to one of the arc magnets' power supply problem.

    08:40 Got call from MCC saying that we may get beam. Moving the target ot BeO to start the long beam recovery procedure since the beam has been down since 2:35 AM.

    08:48 got beam for the spot check.

    09:11 Target moving to 3He - 20 cm.
    09:17 Asked for 1 uA beam for the beam recovery procedure.

    09:21 Asked for 120 uA for production data.

    15:08 Hall C is asking for changing the half wave plate IN->OUT.

    16:00 Beam of the shift, shown in Figure 1, has been very trippy. The blue trace is hall A beam (default value 120 uA) and the gray trace is hall C beam (default value: 165 uA).

    Production run list on 3He 20 cm target at 120 uA:
    4045
    4046
    4047
    4048
    4049
    4050
    4051
    4052
    4053 Half wave plate changed from IN->OUT by Hall C's request
    4054
    Runs 4050-54 had very trippy beam



    FIGURE 1